OpenStudy (anonymous):

find the exact value: cos (45 deg-150 degrees)trying to get this hw done but it timed and i'm running out of time... PLEASE HELP!!

OpenStudy (raden):

use the identity : cos(a-b) =cos(a)cos(b) + sin(a)sin(b) with a = 45 and b = 150

OpenStudy (raden):

cos(45 - 150) = cos(45)cos(150) + sin(45)sin(150) = 1/2 sqrt(2) (-1/2 sqrt(3)) + 1/2 sqrt(2) 1/2 sqrt(2) = - 1/4 sqrt(6) + 1/4 (2) = - 1/4 sqrt(6) + 1/2